A survey was conducted at local colleges around Madison, Wisconsin to find out the average height of a college student. Of 692 students surveyed, 421 replied that they were over 6 feet tall. 


What is the standard error? Answer choices are rounded to the hundredths place. 

a.) 0.24 b.) 0.49 c.) 0.02  d.) 12.84
Answer #1

Here,

Proportion of students over 6 feet \(=\mathrm{p}=421 / 692=0.6084\)

Total number of students \(=\mathrm{n}=692\)

So,

Standard error

\(=\sqrt{\frac{p(1-p)}{n}}=\sqrt{\frac{0.6084 * 0.3916}{692}}\)

\(=0.02\)

So, correct option:

(c) \(0.02\)